OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5

OpenStudy (anonymous):

i will give a medal to anyone who helps

OpenStudy (texaschic101):

The substitution method is just like it sounds...you sub in something for something else. In the equation x = y + 5, we know what x equals. So we sub in y + 5 for x in the other equation. x + y = 6 -- sub in y + 5 for x (y + 5) + y = 6 -- now combine like terms 2y + 5 = 6 -- subtract 5 from both sides 2y = 6 - 5 2y = 1 -- divide both sides by 2 y = 1/2 Now we can sub 1/2 in for y in either equation to find x x + y = 6 x + 1/2 = 6 x = 6 - 1/2 x = 12/2 - 1/2 x = 11/2 now we can check our answers x = y + 5 11/2 = 1/2 + 5 11/2 = 1/2 + 10/2 11/2 = 11/2 (correct) There you go...x = 11/2 and y = 1/2 any questions ?

OpenStudy (anonymous):

x + y = y + 5 (y + 5) + y = 6 x + (y + 5) = 6 x = (x – 6) + 5

OpenStudy (anonymous):

those are the choices